# Predicting and Understanding College Student Mental Health with Interpretable Machine Learning

### v1

@@ Abstract.
Mental health issues among college students have reached critical levels, significantly impacting academic performance and overall wellbeing. Predicting and understanding mental health status among college students is challenging due to three main factors: the necessity for large-scale longitudinal datasets, the prevalence of black-box machine learning models lacking transparency, and the tendency of existing approaches to provide aggregated insights at the population level rather than individualized understanding.
To tackle these challenges, this paper presents I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model, validated on the College Experience Study, the longest longitudinal mobile sensing dataset. This dataset spans five years and captures data from both pre-pandemic periods and the COVID-19 pandemic. I-HOPE connects raw behavioral features to mental health status through five defined behavioral categories as interaction labels. This approach achieves a prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. In addition, our model distills complex patterns into interpretable and individualized insights, enabling the future development of tailored interventions and improving mental health support. The code is available at https://github.com/roycmeghna/I-HOPE.
@@ 1. Introduction
We compare I-HOPE with baseline methods that lack personalized predictions and two-stage feature mappings. I-HOPE achieves an overall prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. I-HOPE simplifies complex patterns into interpretable and personalized insights by mapping behavioral features into interaction labels. For instance, a good 7-hour sleep is linked to better mental health outcomes, while poor sleep correlates with increased anxiety and depression. Walking serves as a stress reliever associated with leisure and relaxation. Phone usage reveals contrasting patterns; interactions at home often indicate positive connections, whereas excessive use in social settings suggests stress or discomfort. The balance between social and personal time is crucial, with shared spaces that promote engagement and personal spaces that allow emotional recharge. Importantly, these behaviors affect individuals differently; some find “Me Time” in working out, while others prefer studying or spending time in their own dorms.
We summarize the contributions as follows:
  • Interpretable machine learning framework development: We present I-HOPE, the first hierarchical model that maps raw behavioral data into five interpretable interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, enhancing the transparency and accuracy of predictions.
@@ 3. Motivation
In this section, we explore the challenges of predicting mental health status by analyzing the dataset and identify key insights for effective solutions.
@@ 3.2. Smartphone Usage Features Alone Are Insufficient for Predicting Mental Health
To assess whether smartphone usage features can effectively predict mental health outcomes, we train a machine learning model on the CES dataset to predict the PHQ-4 score. Given that the PHQ-4 score has four categories, as illustrated in Table 1, this constitutes a multi-class classification problem. Therefore, we use two smartphone features mentioned above—the number of unlocks and daily phone usage duration—as input features to train a multilayer perceptron (MLP) model, a neural network widely used for classification and prediction (Popescu et al., 2009). We preprocess the data by normalizing the features and ensuring balanced class distributions through oversampling for the PHQ-4 categories. We randomly split the dataset into 80% for training and 20% for testing. Our results reveal that the prediction accuracy is only 28%, which is low. Even after oversampling to achieve balanced class distributions for the PHQ-4 categories, the accuracy rises only to 30%. These results indicate that relying solely on these smartphone usage features is insufficient to predict mental health outcomes accurately.
@@ 3.3. Improving Predictions with Additional Features
To improve prediction accuracy, we need to incorporate additional features from the CES dataset. Specifically, we add 33 more features from the CES dataset related to location information, sleep duration, and activity times (see Column 2 of Table 3). With increased input dimensions, we use the same MLP model with 35 input nodes, three fully connected hidden layers, and four output nodes corresponding to the PHQ-4 categories. The training and test sets are the same as above. The model is trained for 50 epochs until the learning curve stabilizes. This approach improves the prediction accuracy to 60%, which is still not high. We attribute the low accuracy to two reasons: (i) Treating the dataset as a whole may have obscured individual differences; (ii) A high correlation among features likely reduce the model’s ability to make accurate predictions. We discuss these below:
@@
The CES dataset includes data from 217 individuals, each contributing about 160 data points. Research indicates that individual differences in behavior and preferences can significantly affect predictive modeling outcomes (He et al., 2023; Lengerich et al., 2019). To capture these variations, we develop personalized models by grouping data based on each student’s unique user ID (UID). Each model uses the same 35 features but is trained on the corresponding individual’s data, maintaining the original neural network architecture and train-test split. This personalized approach improves the average prediction accuracy to 70%. However, high feature correlation remains a challenge, introducing redundancy and limiting further accuracy improvements, highlighting the need for additional refinement to address these feature-related issues effectively.
@@
Highly correlated features introduce redundancy, leading to overfitting and reduced test accuracy while increasing computational complexity without adding valuable information. As shown in Figure 2, many dataset features are highly correlated. To address this, we analyze feature importance using random forest, an efficient and interpretable ensemble learning technique (Breiman, 2001). This method builds multiple decision trees and combines their outputs to evaluate feature importance. Figure 3 shows the importance ranking of all 35 features, highlighting the top ones. We select the top 50% of these features to retrain personalized neural network models, adjusting the input dimensions accordingly. However, this results in a drop in prediction accuracy to 65%, indicating that simply excluding certain features may remove critical information necessary for effective predictions.
@@
These analyses suggest that the features influencing prediction accuracy vary among individuals. For instance, a higher biking time may correlate with better mental health for one person, while walking might be more relevant for another, and a combination of running and walking could be key for someone else. A global feature importance approach risks eliminating crucial features for individuals, potentially reducing accuracy. Therefore, it is vital to incorporate personalization while minimizing feature correlation, ensuring the model captures individual-specific patterns without being hindered by redundant or irrelevant features, ultimately enhancing overall prediction accuracy.
@@ 4. I-HOPE
So far, our analysis has indicated that globally selecting features for model training reduces correlation but does not enhance prediction accuracy. This underscores the need for a personalized approach that minimizes feature correlation while capturing individual-specific patterns, improving prediction accuracy and interpretability.
Therefore, we design and implement I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects behavioral features to mental health status through an intermediate layer of five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, which categorize different daily behaviors. This structure offers two benefits:
@@
    Enhanced Interpretability. I-HOPE enables the identification of an individual’s social type and highlights the interactions that influence their mental health the most.
Figure 4 outlines the workflow of I-HOPE design. It starts with n raw input features ($n=35$ from the CES dataset), which are mapped to k interaction labels ($k=5$), representing aggregated behavioral categories: Leisure, Me Time, Phone Time, Sleep, and Social Time. To quantify the behavioral insights, we define a score for each label and use the scores as input for Stage 2. I-HOPE uses a thresholding mechanism that incorporates personalized feature importance to calculate interaction label scores, effectively reducing feature correlation by condensing the 35 raw features into five labels. These scores serve as input to a neural network that predicts one of four mental health categories based on the PHQ-4 scale: Normal, Mild, Moderate, or Severe. I-HOPE groups individual preferences under interaction labels, improving predictive consistency. It also improves interpretability by identifying which interaction label score influences the predictions and determining the significance of specific raw features within each label. For example, I-HOPE can assess whether “walking” or “spending time with friends” in the Leisure category has a greater impact on mental health of an individual. Next, we will describe each stage of I-HOPE and explain the threshold mechanism that links them.
@@ 4.1. Stage 1: Input Features $\rightarrow$ Interaction Labels
In Stage 1, I-HOPE maps raw input features to five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time. These labels group features with statistically significant associations to mental health outcomes (p¡0.05 or p¡0.01) into each label. This approach combines domain knowledge and data-driven insights to capture relevant behavioral dimensions, reduce redundancy, and enhance interpretability while maintaining granularity for personalized analysis. The mapping process is described below.
@@
    Feature selection and initial mapping: We manually select a subset of features from the dataset ($n=35$) to match each interaction label based on their semantic and behavioral relevance. For instance, features like unlock duration and phone conversations belong to Phone Time, while duration of social activities and time spent with friends belong to Social Time. The initial mapping is detailed in Table 3. Notably, some features are assigned to multiple labels, reflecting their multidimensional impact on various behavioral categories.
@@
    Feature clustering: To group features with similar behavioral patterns and ensure consistency within interaction labels, I-HOPE uses K-means clustering to identify inherent groupings that capture meaningful behavioral dimensions without predefined ground truth.
@@
    Feature importance analysis: I-HOPE uses a random forest model to assess the contribution of individual input features to each interaction label, as it can handle high-dimensional data and capture nonlinear relationships (Breiman, 2001). Random forest measures feature importance by evaluating how much each feature reduces uncertainty at decision splits across all trees. Features that significantly improve data separation receive higher importance scores, which are then used to refine interaction label scoring. This analysis is conducted per-user basis, grouping data by UIDs.
Algorithm 1 Calculation of LeisureScore
1:  Input: Threshold of features mapped to Leisure across UIDs , Feature Importances (FeatImps)
2:  Initialize: LeisureScore = 0
3:  if Walking Duration $<$ Threshold (where Threshold = Mean of distribution) then
4:     LeisureScore += FeatImp of “Walking Duration”
@@
6:  if Phone usage at home $<$Threshold (where Threshold = Mean of distribution) then
7:     LeisureScore += FeatImp of “Phone usage at home”
@@
9:  if Running Duration $<$ Threshold (where Threshold = Mean of distribution) then
10:     LeisureScore += FeatImp of “Running Duration”
@@
12:  if Duration of Conversation $>$ Threshold (where Threshold = Mean of distribution) then
13:     LeisureScore += FeatImp of “Duration of Conversation”
@@
15:  if Phone usage at others’ dorm $<$ Threshold (where Threshold = Mean of distribution) then
16:     LeisureScore += FeatImp of “Phone usage at others’ dorm”
17:  end if
18:  if Voice Detection in others’ dorm $>$ Threshold (where Threshold = Mean of distribution) then
19:     LeisureScore += FeatImp of “Voice Detection in others’ dorm”
20:  end if
21:  if Duration of conversation in others’ dorms $>$ Threshold (where Threshold = Mean of distribution) then
22:     LeisureScore += FeatImp of Duration of conversation in others’ dorms
23:  end if
Algorithm 2 Calculation of SleepScore
1:  Input: Threshold of features mapped to label Sleep across UIDs, Feature Importances (FeatImps)
2:  Initialize: SleepScore = 0
3:  if Sleep duration $>$ Threshold (where Threshold = Mean of distribution) then
4:     SleepScore += FeatImp of “Sleep duration”
5:  end if
6:  if Duration at own dorm $>$ Threshold (where Threshold = Mean of distribution) then
7:     SleepScore += FeatImp of “Duration at own dorm”
8:  end if
9:  if Duration of stillness $>$ Threshold (where Threshold = Mean of distribution) then
10:     SleepScore += FeatImp of “Duration of stillness”
11:  end if
12:  if Conversation detected at home $<$ Threshold (where Threshold = Mean of distribution) then
13:     SleepScore += FeatImp of “Conversation detected at home”
14:  end if
4.2. Threshold based Scoring
This section outlines the scoring mechanism for interaction labels. Each label captures relevant behavioral information, allowing us to assess the importance of features for individuals. Since these labels are qualitative, they must be converted into quantitative features for predictive analysis. To achieve this, we define a score, termed [InteractionLabel]Score, for each label based on thresholds derived from the distribution of associated characteristics. The score is updated by checking if a feature’s value exceeds or falls below the threshold and adjusting it according to its importance.
To ensure that the scores are tailored to individual users, I-HOPE updates the interaction label scores by assigning greater weight to features with higher importance scores. This personalized approach enables a more accurate representation of each label and enhances the relevance of the features for predictive modeling. Each label’s score is the sum of feature contributions that meet the threshold, where contributions are weighted based on feature importance. These scores are then used as input features for predictive modeling.
To demonstrate this methodology, we use the labels Leisure and Sleep as examples. These labels are selected for clarity and simplicity, but the same process applies uniformly to all other labels.
  • The LeisureScore is derived using the following features (Algorithm 1): duration of walking, ratio of phone unlocks to duration of phone use, duration of running, duration of conversation, ratio of phone unlocks to duration of phone use in others’ dorms, and duration of conversation in others’ dorms.
  • The SleepScore is derived using the following features (Algorithm 2): sleep duration, duration of the phone being idle, time spent in own dorm, and duration of conversation detected.
@@ 4.3. Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Category
In Stage 2, I-HOPE uses the five interaction label scores (LeisureScore, MeScore, PhoneScore, SleepScore, SocialScore) as inputs to predict PHQ-4 categories. This stage simplifies the predictive process by using these labels as compact and interpretable representations of the raw data. The process includes the following steps:
  1. (1)
    Neural network prediction: I-HOPE builds a neural network model to predict PHQ-4 categories using the five interaction label scores as inputs. The neural network includes an input layer with five nodes, three hidden layers for pattern extraction, and an output layer with four nodes for the PHQ-4 categories (normal, mild, moderate, and severe). I-HOPE uses the ReLU activation function for hidden layers and the Softmax activation function for the output layer for multi-class classification. The neural network model is trained for 50 epochs using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 and categorical cross-entropy as the loss function.
  2. (2)
    Weighted feature representation: The input scores of the neural network are weighted by the importance of the feature of the random forest analysis, ensuring that the model captures individual variations and enhances the accuracy of prediction.
  3. (3)
    Performance evaluation: I-HOPE calculates the mean test accuracy across all UIDs to measure performance while accounting for individual differences. Using interaction labels reduces input complexity, highlighting the effectiveness of I-HOPE’s hierarchical design.

### v2

@@ Abstract.
Mental health issues among college students have reached critical levels, affecting both academic performance and overall wellbeing. Predicting and understanding mental health status among college students is challenging due to three key barriers: the lack of large-scale longitudinal datasets, the prevalence of black-box machine learning models that offer little transparency, and a reliance on population-level analysis rather than personalized understanding.
To tackle these challenges, this paper presents I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects raw behavioral features to mental health status through five defined behavioral categories as interaction labels. We evaluate I-HOPE on the College Experience Study, the longest longitudinal mobile sensing dataset. This dataset spans five years and captures data from both pre-pandemic periods and the COVID-19 pandemic. I-HOPE achieves a prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. In addition, I-HOPE distills complex patterns into interpretable and individualized insights, enabling the future development of tailored interventions and improving mental health support. The code is available at https://github.com/roycmeghna/I-HOPE.
@@ 1. Introduction
We compare I-HOPE with baseline methods that lack personalized predictions or two-stage feature mappings. I-HOPE achieves an overall prediction accuracy of 91%, significantly surpassing the 60-70% accuracy of baseline methods. I-HOPE simplifies complex patterns into interpretable and personalized insights by mapping behavioral features into interaction labels. For instance, a good 7-hour sleep is linked to better mental health outcomes, while poor sleep correlates with increased anxiety and depression. Walking serves as a stress reliever associated with leisure and relaxation. Phone usage reveals contrasting patterns; interactions at home often indicate positive connections, whereas excessive use in social settings suggests stress or discomfort. The balance between social and personal time is crucial, with shared spaces that promote engagement and personal spaces that allow emotional recharge. Importantly, these behaviors affect individuals differently; some find “Me Time” in working out, while others prefer studying or spending time in their own dorms.
We summarize the contributions as follows.
  • Interpretable machine learning framework development: We present I-HOPE, the first hierarchical model that maps raw behavioral data into five interpretable interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, enhancing the transparency and prediction accuracy.
@@ 3. Motivation
In this section, we explore the challenges of predicting mental health status by analyzing the CES dataset and identify key insights for effective solutions.
@@ 3.2. Smartphone Usage Features Alone Are Insufficient for Predicting Mental Health
To assess whether smartphone usage features can effectively predict mental health outcomes, we train a machine learning model on the CES dataset to predict the PHQ-4 score. Given that the PHQ-4 score has four categories, as illustrated in Table 1, this constitutes a multi-class classification problem. Therefore, we use two smartphone features mentioned above—the number of unlocks and daily phone usage duration—as input features to train a multilayer perceptron (MLP) model, a neural network widely used for classification and prediction (Popescu et al., 2009). The MLP model has 2 input nodes, 3 fully connected hidden layers, and 4 output nodes corresponding to the PHQ-4 categories. The model is trained for 50 epochs until the learning curve stabilizes. We preprocess the data by normalizing the features and ensuring balanced class distributions through oversampling for the PHQ-4 categories. We randomly split the dataset into 80% for training and 20% for testing. Our results reveal that the prediction accuracy is only 28%, which is low. Even after oversampling to achieve balanced class distributions for the PHQ-4 categories, the accuracy rises only to 30%. These results indicate that relying solely on these smartphone usage features is insufficient to predict mental health outcomes accurately.
@@ 3.3. Improving Predictions with Additional Features
To improve prediction accuracy, we need to incorporate additional features from the CES dataset. Specifically, we use 35 mannully selected features from the CES dataset related to location information, sleep duration, and activity times (see Column 2 of Table 3). We use the MLP model with 35 input nodes, 3 fully connected hidden layers, and 4 output nodes corresponding to the PHQ-4 categories. The training and test sets are the same as above. The model is trained for 50 epochs until the learning curve stabilizes. This approach improves the prediction accuracy to 60%, which is still not high. We attribute the low accuracy to two reasons: (i) treating the dataset as a whole may have obscured individual differences; (ii) a high correlation among features likely reduce the model’s ability to make accurate predictions. We discuss these below.
@@
The CES dataset includes data from 217 individuals, each contributing about 160 data points. Research indicates that individual differences in behavior and preferences can significantly affect predictive modeling outcomes (He et al., 2023; Lengerich et al., 2019). To capture these variations, we develop personalized MLP models by grouping data based on each student’s unique user ID (UID). Each model uses the same 35 features but is trained on the corresponding individual’s data, maintaining the original MLP architecture and train-test split. This personalized approach improves the average prediction accuracy to 70%. However, high feature correlation remains a challenge, introducing redundancy and limiting further accuracy improvements, highlighting the need for additional refinement to address these feature-related issues effectively.
@@
Highly correlated features introduce redundancy, leading to overfitting and reduced test accuracy while increasing computational complexity. As shown in Figure 2, many features are highly correlated. To address this, we analyze feature importance using random forest, an efficient and interpretable ensemble learning technique (Breiman, 2001). This method builds multiple decision trees and combines their outputs to evaluate feature importance. Figure 3 shows the importance ranking of all 35 features, highlighting the top ones. We select the top 50% of these features to retrain personalized MLP models, adjusting the input dimensions accordingly. However, this results in a drop in prediction accuracy to 65%, indicating that simply excluding certain features may remove critical information necessary for accurate predictions.
@@
These analyses suggest that the features influencing prediction accuracy vary among individuals. For instance, a higher biking time may correlate with better mental health for one person, while walking may be more relevant for another, and a combination of running and walking could be key for someone else. A global feature importance approach risks eliminating crucial features for individuals, potentially reducing accuracy. Therefore, it is vital to incorporate personalization while minimizing feature correlation, ensuring the model captures individual-specific patterns without being hindered by redundant or irrelevant features, ultimately enhancing overall prediction accuracy.
@@ 4. I-HOPE
Our analysis has shown that globally selecting features for model training reduces correlation but does not enhance prediction accuracy. This highlights the need for a personalized approach that reduces feature correlation and captures individual-specific patterns, leading to better prediction accuracy and interpretability.
Therefore, we design I-HOPE, the first Interpretable Hierarchical mOdel for Personalized mEntal health prediction. I-HOPE is a two-stage hierarchical model that connects behavioral features to mental health status through an intermediate layer of five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time, which categorize different daily behaviors. This model structure offers two benefits:
@@
    Enhanced Interpretability. I-HOPE enables the identification of an individual’s social type and highlights the interactions that correlate their mental health the most.
Figure 4 presents the workflow of I-HOPE design. It starts with n raw input features ($n=35$ from the CES dataset), which are mapped to k interaction labels ($k=5$), representing aggregated behavioral categories: Leisure, Me Time, Phone Time, Sleep, and Social Time. To quantify the behavioral insights, we assign a score for each interaction label by introducing a two-step rule-based method. These scores then serve as input to a neural network that predicts one of four mental health categories: Normal, Mild, Moderate, or Severe. In summary, by leveraging the hierarchical model structure, I-HOPE improves both prediction accuracy and model interpretability. For example, I-HOPE can assess whether “walking” or “spending time with friends” in the Leisure category has a greater impact on mental health of an individual. I-HOPE is conducted per-user basis, grouping data by user IDs. Next, we will describe each stage of I-HOPE in detail.
@@ 4.1. Stage 1: Input Features $\rightarrow$ Interaction Labels
In Stage 1, I-HOPE maps raw input features to five interaction labels: Leisure, Me Time, Phone Time, Sleep, and Social Time. These labels group features with statistically significant associations to mental health outcomes (p¡0.05 or p¡0.01) into each label. This approach combines domain knowledge and data-driven insights to capture relevant behavioral dimensions, reduce redundancy, and enhance interpretability while maintaining granularity for personalized analysis. The detailed process is described below.
Algorithm 1 Rule-based initialization of SleepScore
1:  Input: Thresholds of each feature in Sleep label, which are population means of the feature distributions.
2:  Initialize: SleepScore = 0
3:  if Sleep duration $>$ Threshold (where Threshold = Mean of Sleep duration distribution) then
4:     SleepScore += 1
5:  end if
6:  if Duration at own dorm $>$ Threshold (where Threshold = Mean of Duration at own dorm distribution) then
7:     SleepScore += 1
8:  end if
9:  if Duration of stillness $>$ Threshold (where Threshold = Mean of Duration of stillness distribution) then
10:     SleepScore += 1
11:  end if
12:  if Conversation detected at home $<$ Threshold (where Threshold = Mean of Conversation detected at home distribution) then
13:     SleepScore += 1
14:  end if
15:  Output: SleepScore
@@
    Feature selection and initial mapping: We manually select a subset of features from the dataset ($n=35$) to match each interaction label based on their semantic and behavioral relevance. For instance, features like unlock duration and phone conversations belong to Phone Time, while duration of social activities and time spent with friends belong to Social Time. The initial mapping is detailed in Table 3. Notably, some features are assigned to multiple labels, reflecting their multidimensional impact on various behavioral categories. We further employ K-means clustering with k=5 to perform a soft validation, assessing whether five interaction labels are a reasonable choice.
@@
    Rule-based scoring initialization: Since interaction labels are manually created, we need to assign them numerical values for downstreaming quantitative analysis. I-HOPE uses a two-step rule-based scoring method to compute an [InteractionLabel]Score for each label. The first step is to initialize the scores. Specifically, for each feature within a label, we compare its value to the population mean value of that feature; if it exceeds the mean, the score increases by 1. The intuition is that higher feature values suggest stronger relevance to the label and thus contribute more to its score. We illustrate the first step of this method using the Sleep label in Algorithm 1, though the same procedure applies to all interaction labels.
@@
    Rule-based scoring using feature importance: The second step of the rule-based scoring method is to use a random forest model on the features and initialized scores to assess the contribution of individual input features to each interaction label. We choose random forest due to its ability to handle high-dimensional data and capture nonlinear relationships (Breiman, 2001). I-HOPE then uses feature importance generated from random forest to refine the final scores. Specifically, for each feature within a label, we compare its value to the population mean; if it exceeds the mean, the score increases by its corresponding feature importance value. We illustrate the second step of this method using the Sleep label in Algorithm 2, though the same procedure applies to all interaction labels.
Algorithm 2 Rule-based final scoring of SleepScore
1:  Input: Thresholds of each feature in Sleep label, which are population means of the feature distributions.
2:  Initialize: SleepScore = 0
3:  if Sleep duration $>$ Threshold (where Threshold = Mean of Sleep duration distribution) then
4:     SleepScore += FeatImp of Sleep duration
@@
6:  if Duration at own dorm $>$ Threshold (where Threshold = Mean of Duration at own dorm distribution) then
7:     SleepScore += FeatImp of Duration at own dorm
@@
9:  if Duration of stillness $>$ Threshold (where Threshold = Mean of Duration of stillness distribution) then
10:     SleepScore += FeatImp of Duration of stillness
@@
12:  if Conversation detected at home $<$ Threshold (where Threshold = Mean of Conversation detected at home distribution) then
13:     SleepScore += FeatImp of Conversation detected at home
@@
15:  Output: SleepScore
@@ 4.2. Stage 2: Interaction Labels $\rightarrow$ PHQ-4 Category
In Stage 2, I-HOPE uses the five interaction label scores (LeisureScore, MeScore, PhoneScore, SleepScore, SocialScore) as inputs to predict PHQ-4 categories. This stage simplifies the predictive process by using these labels as compact and interpretable representations of the raw data. Specifically, I-HOPE builds a neural network model to predict PHQ-4 categories using the five interaction label scores as inputs. The neural network includes an input layer with five nodes, three hidden layers for pattern extraction, and an output layer with four nodes for the PHQ-4 categories (Normal, Mild, Moderate, or Severe). I-HOPE uses the ReLU activation function for hidden layers and the Softmax activation function for the output layer for multi-class classification. The neural network model is trained for 50 epochs using the Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 and categorical cross-entropy as the loss function.